OpenStudy (anonymous):

Use the given info to determine the exact trigonometric value. tanX = (square root of 3)/4, 0

zepdrix (zepdrix):

Recall:\[\Large \cot x=\frac{1}{\tan x}\]

OpenStudy (anonymous):

Thanks, I have no idea what this given info means, specifically 0<X<(pi/2). Could you maybe help me understand?

OpenStudy (anonymous):

I get what it means, but not how to use it basically

zepdrix (zepdrix):

It doesn't actually matter in this particular case. But in general it's telling us `which quadrant` our angle X falls into. Since our angle is between 0 and pi/2, we're in the first quadrant, yes?

zepdrix (zepdrix):

With tangent and cotangent, they will always share the same `sign`, so we don't have to worry about making a mistake with this particular problem. (tangent is positive, so cotangent will be positive).

zepdrix (zepdrix):

So like ummm...

zepdrix (zepdrix):

Here's a good example problem for you:

zepdrix (zepdrix):

Use the given info to determine the exact trigonometric value. tanX = 3 / 4 ; pi < X < (3pi/2) ; sinX = ?

zepdrix (zepdrix):

Understand how the `pi < X < (3pi/2)` is telling us that we're in the 3rd quadrant?

OpenStudy (anonymous):

Wow okay, yeah, that really helped. Thank you!